# Efficient Sampling for Pauli Measurement-Based Shadow Tomography in Direct Fidelity Estimation

### v6

@@ I Introduction
Advances in the preparation of many-body entangled quantum states have marked significant progress in the field of quantum information [1, 2, 3, 4]. A critical step in these experiments is verifying that the system’s state aligns with the desired one [5, 6, 7, 8]. Traditional methods for assessing quantum state fidelity include quantum state tomography (QST), which reconstructs the full density matrix of a quantum state, offering detailed information [9, 10, 11, 12, 13, 14, 15, 16, 17]. However, QST is resource-intensive and scales poorly with system size, posing significant challenges for large-scale quantum systems [18]. Recent approaches have introduced various techniques aimed at simplifying the processes of verification [19, 20, 21, 22] and fidelity estimation [23, 24, 25]. Among these, direct fidelity estimation (DFE) has gained attention for its potential to provide accurate fidelity measurements with fewer resources compared to full state tomography. Unlike full tomography, DFE focuses on directly estimating the fidelity without requiring a complete reconstruction of the quantum state, thus significantly reducing the measurement overhead. For some classes of states, efficient methods using Pauli measurement [26, 23, 24, 27], some relying on importance sampling [28, 29], are known.
@@ III Classical shadows
    Upon receiving the measurement outcome $\hat{b}\in\{0,1\}^{n}$, store an efficient description of $U^{\dagger}|\hat{b}\rangle\langle\hat{b}|U$ in classical memory.
@@ IV.1 Random $n$-qubit Clifford circuits
where $\mathbb{I}$ denotes the $2^{n}$-dimensional identity operator. A classical shadow of size independent of $n$ is sufficient to accurately estimate the fidelity with any pure target state [30, Theorem 1].
@@ IV.2 Tensor products of random single-qubit Clifford circuits
The worst case upper bound on the number of measurements to estimate (3) scales exponentially with the locality of $O$ [30, Proposition 3] or the stabilizer Rényi entropy [27, Theorem 1]. However, there exist much efficient methods for DFE with some well-conditioned states (as defined in Ref. [23]), despite the locality of $O$ being $n$.
@@ V.1 Fidelity with computational basis states
Note that $|\langle\hat{b}_{i}|U_{i}|b_{i}\rangle|=1/\sqrt{2}$ whenever $U_{i}\in T\setminus\{\mathbb{I}\}$, which can be verified with simple calculations. That is, if we measure the $i$th qubit in the $X$ or $Y$ basis, then $13|\langle\hat{b}_{i}|U_{i}|b_{i}\rangle|^{2}-$ 1 is always $1/2$ (regardless of $\hat{b}_{i}$). This allows us to obtain a sample for each of the $3^{n}$ Pauli measurements just from a single $ZZ\cdots Z$ measurement outcome $|\hat{b}\rangle$, because we can replace some of the $Z$ measurements with $X$ or $Y$ measurements for which the multiplier is $1/2$ with certainty. Let $m=|\{i|\hat{b}_{i}=b_{i}\}|$. Substituting $U_{i}$ with $\mathbb{I}$ for all $i$ in (V.1) gives
@@
$\langle 0|U_{i}^{\dagger}|0\rangle\langle 0|U_{i}|1\rangle$ $1/2$ $-i/2$ 00
$\langle 0|U_{i}^{\dagger}|1\rangle\langle 1|U_{i}|1\rangle$ $-1/2$ $i/2$ 00
$\langle 1|U_{i}^{\dagger}|0\rangle\langle 0|U_{i}|0\rangle$ $1/2$ $i/2$ 00
$\langle 1|U_{i}^{\dagger}|1\rangle\langle 1|U_{i}|0\rangle$ $-1/2$ $-i/2$ 00
@@ V.2 Fidelity with GHZ state
evaluates to ±$\pm 1/2^{n-1}$ if (9) holds and zero otherwise. This symmetry allows us to further reduce the sample space of $(U_{i})$ as
@@
[…]
@@
    Sample $U_{i}\in\{H,HS^{\dagger}\}$ uniformly and independently for $11\leq i\leq n-$ 1.
    Choose $U_{n}\in\{H,HS^{\dagger}\}$ such that |$|\{i\mid U_{i}=HS^{\dagger}\}|\mod{2}=0$.
@@
where $\hat{o}\equiv 2^{n-2}V$. Therefore, we have $20\leq\hat{d}\leq 1/$ 2 and $-1/2\leq\hat{o}\leq 1/2$.
Finally, $\hat{F}$ is defined as a sample generated by sampling $3(\hat{d}-1/4)$ with probability $1/3$ and $23\hat{o}/$ 2 with probability $2/3$. Then $|\hat{F}|\leq 3/4$ and
@@
The total number of measurements is about $9\log(2/\delta)/8\epsilon^{2}$, which tightens the bound (2) by a factor of $9128(1+\log_{2/\delta}2)/$ 9, a constant between $128/9$ and $256/9$ (closer to $128/9$ for small $\delta$). Alg. 1 provides a complete pseudocode for the GHZ state. It should be noted that the measurement settings for fidelity estimation in Alg. 1 also appear in some verification protocols [53, 54], but in our context these settings are further associated with a fidelity estimator defined by the specific measurement basis and outcome.
@@ V.3 Fidelity with W state
The off-diagonal estimator $\hat{o}$ can be found in Appendix A. We have $0\leq\hat{d}\leq 1/n$ and $-(n-1)/2\leq\hat{o}\leq(n-1)/2$.
@@
[…]
@@ V.4 Fidelity with Dicke state
We have $0\leq\hat{d}\leq\binom{n}{k}^{-1}$ and $-\binom{n}{k}^{-1}\cdot c_{l}\leq\hat{o}_{l}\leq\binom{n}{k}^{-1}\cdot c_{l}$, where the definition of $c_{l}$ and the off-diagonal estimator $\hat{o}_{l}$ can be found in Appendix B. Moreover, let
@@
[…]
@@
            + |$|\{m\in\text{s}(\mathbf{j})-\text{s}(\mathbf{i})\mid\hat{b}_{m}=1\}|$
[…]
@@ VI A generic optimization rule
After restricting the sampling domain to (16) we reweight and renormalize the estimators and the sampling probabilities (which is itself a nontrivial process that depends on both the set (16) and the coefficients $c_{b_{1},b_{2}}$), replacing the original shadow estimation using all $3^{n}$ Pauli operators. An immediate corollary is that Pauli strings containing an odd number of $Y$ are never required when $O$ is real.
@@ Appendix A Off-diagonal terms for W state
where $[n]\equiv\{1,2,\cdots,n\}$. The third equality follows from previous analyses to reduce the sample space.
From Table 1, we see that the term inside the square bracket in (A) is zero if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=0$, ±$\pm 1/4$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}=U_{j}$, and ±$\pm i/4$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}\neq U_{j}$. Meanwhile, $\text{tr}(\rho|e_{j}\rangle\langle e_{i}|)$ is the complex conjugate of (A), and the sample
@@
evaluates to ±$\pm 1/2$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}=U_{j}$. Otherwise, it is always zero. This symmetry allows us to further reduce the sample space of $(U_{i})$:
@@
where $\hat{o}\equiv(n-1)V_{\text{W}}$. Therefore, we have $-(n-1)/2\leq\hat{o}\leq(n-1)/2$.
@@ Appendix B Off-diagonal terms for Dicke state
$\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}\right],$ (19)
@@
From Table 1, we see that the term inside the square bracket in (B) is zero if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}=0$ or $\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=0$, ±$\pm 1/2^{2k-2l}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and |$|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 0\pmod{2}$, and ±$\pm i/2^{2k-2l}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and |$|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 1\pmod{2}$. Meanwhile, $\text{tr}(\rho|\mathbf{j}\rangle\langle\mathbf{i}|)$ is the complex conjugate of (B), and the sample
@@
evaluates to ±$\pm 1/2^{2k-2l-1}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and |$|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 0\pmod{2}$. Otherwise, it is always zero. This symmetry allows us to further reduce the sample space of $(U_{m})$:
@@
$o_{l}$ $=\sum_{\begin{subarray}{c}\mathbf{i}<\mathbf{j},\\ |\mathbf{i}|=|\mathbf{j}|=k,\\ |\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})|=l\end{subarray}}\frac{1}{c_{l}}\sum_{U_{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})-\max(\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j}))}^{\prime}\in T\setminus\{\mathbb{I}\}}\frac{1}{2^{2k-2l-1}}\mathbb{E}_{\hat{b}}\left[\hat{o}_{l}\left|\begin{aligned} U_{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})-\max(\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j}))}=U_{m}^{\prime},\\ U_{\max(\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j}))}=\mathcal{F}(\{U_{m}^{\prime}\}),\\ U_{m\notin\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})}=\mathbb{I}\end{aligned}\right.\right]$
$=\mathbb{E}_{\{\mathbf{i},\mathbf{j}\}\sim f_{l},U_{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})-\max(\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j}))}^{\prime}\sim g,\hat{b}}\left[\hat{o}_{l}\left|\begin{aligned} U_{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})-\max(\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j}))}=U_{m}^{\prime},\\ U_{\max(\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j}))}=\mathcal{F}(\{U_{m}^{\prime}\}),\\ U_{m\notin\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})}=\mathbb{I}\end{aligned}\right.\right],$
where $\hat{o}_{l}\equiv V_{\text{D}}\cdot\binom{n}{k}^{-1}\cdot c_{l}\cdot 2^{2k-2l-1}$ and ($\mathcal{F}(\{U_{m}^{\prime}\})\in T\setminus\{\mathbb{I}\}$ such that |$|\{U\in\{U_{m}^{\prime}\}\cup\mathcal{F}(\{U_{m}^{\prime}\})|U=HS^{\dagger}\}|\equiv 0\pmod{2}$. Therefore, we have $-\binom{n}{k}^{-1}\cdot c_{l}\leq\hat{o}_{l}\leq\binom{n}{k}^{-1}\cdot c_{l}$.

### v8

# Efficient sampling for Pauli-measurement-based shadow tomography in direct fidelity estimation

@@ I Introduction
Advances in the preparation of many-body entangled quantum states have marked significant progress in the field of quantum information [1, 2, 3, 4]. A critical step in these experiments is verifying that the system’s state aligns with the desired one [5, 6, 7, 8]. One application of this verification step is in the verifier-prover paradigm of cloud quantum computing [9, 10]. Here, an untrusted server (prover) prepares an entangled many-body state, and the client (verifier) must certify that the returned state matches the target. This is accomplished by the prover supplying multiple copies of the state, which the verifier then measures to validate the preparation. Traditional methods for assessing quantum state fidelity include quantum state tomography (QST), which reconstructs the full density matrix of a quantum state, offering detailed information [11, 12, 13, 14, 15, 16, 17, 18, 19]. However, QST is resource-intensive and scales poorly with system size, posing significant challenges for large-scale quantum systems [20]. Recent approaches have introduced various techniques aimed at simplifying the processes of verification [10, 21, 22, 23] and fidelity estimation [24, 25, 26]. Among these, direct fidelity estimation (DFE) has gained attention for its potential to provide accurate fidelity measurements with fewer resources compared to full state tomography. Unlike full tomography, DFE focuses on directly estimating the fidelity without requiring a complete reconstruction of the quantum state, thus significantly reducing the measurement overhead. For some classes of states, efficient methods using Pauli measurement [27, 24, 25, 28], some relying on importance sampling [29, 30], are known.
@@ III Classical shadows
[…]
@@ IV.1 Random $n$-qubit Clifford circuits
where $\mathbb{I}$ denotes the $2^{n}$-dimensional identity operator. A classical shadow of size independent of $n$ is sufficient to accurately estimate the fidelity with any pure target state [31, Theorem 1].
@@ IV.2 Tensor products of random single-qubit Clifford circuits
The worst case upper bound on the number of measurements to estimate (3) scales exponentially with the locality of $O$ [31, Proposition 3] or the stabilizer Rényi entropy [28, Theorem 1]. However, there exist much efficient methods for DFE with some well-conditioned states (as defined in Ref. [24]), despite the locality of $O$ being $n$.
@@ V.1 Fidelity with computational basis states
Note that $|\langle\hat{b}_{i}|U_{i}|b_{i}\rangle|=1/\sqrt{2}$ whenever $U_{i}\in T\setminus\{\mathbb{I}\}$, which can be verified with simple calculations. That is, if we measure the $i$th qubit in the $X$ or $Y$ basis, then $3|\langle\hat{b}_{i}|U_{i}|b_{i}\rangle|^{2}-1$ is always $1/2$ (regardless of $\hat{b}_{i}$). This allows us to obtain a sample for each of the $3^{n}$ Pauli measurements just from a single $ZZ\cdots Z$ measurement outcome $|\hat{b}\rangle$, because we can replace some of the $Z$ measurements with $X$ or $Y$ measurements for which the multiplier is $1/2$ with certainty. Let $m=|\{i|\hat{b}_{i}=b_{i}\}|$. Substituting $U_{i}$ with $\mathbb{I}$ for all $i$ in (V.1) gives
@@
$\langle 0|U_{i}^{\dagger}|0\rangle\langle 0|U_{i}|1\rangle$ $1/2$ i/2-i/2 0
$\langle 0|U_{i}^{\dagger}|1\rangle\langle 1|U_{i}|1\rangle$ 1/2-1/2 $i/2$ 0
$\langle 1|U_{i}^{\dagger}|0\rangle\langle 0|U_{i}|0\rangle$ $1/2$ $i/2$ 0
$\langle 1|U_{i}^{\dagger}|1\rangle\langle 1|U_{i}|0\rangle$ 1/2-1/2 i/2-i/2 0
@@ V.2 Fidelity with GHZ state
evaluates to $\pm 1/2^{n-1}$ if (9) holds and zero otherwise. This symmetry allows us to further reduce the sample space of $(U_{i})$ as
@@
    $\texttt{sum}\,\texttt{+=}\,3\left(\delta_{\hat{b},\mathbf{0}}+\delta_{\hat{b},\mathbf{1}}\right)/2-3/4$
@@
    Sample $U_{i}\in\{H,HS^{\dagger}\}$ uniformly and independently for $1\leq i\leq n-1$.
    Choose $U_{n}\in\{H,HS^{\dagger}\}$ such that $|\{i\mid U_{i}=HS^{\dagger}\}|\mod{2}=0$.
@@
where $\hat{o}\equiv 2^{n-2}V$. Therefore, we have $0\leq\hat{d}\leq 1/2$ and $1/2\leq\hat{o}\leq 1/2$.
Finally, $\hat{F}$ is defined as a sample generated by sampling $3(\hat{d}-1/4)$ with probability $1/3$ and $3\hat{o}/2$ with probability $2/3$. Then $|\hat{F}|\leq 3/4$ and
@@
The total number of measurements is about $9\log(2/\delta)/8\epsilon^{2}$, which tightens the bound (2) by a factor of $128(1+\log_{2/\delta}2)/9$, a constant between $128/9$ and $256/9$ (closer to $128/9$ for small $\delta$). Alg. 1 provides a complete pseudocode for the GHZ state. It should be noted that the measurement settings for fidelity estimation in Alg. 1 also appear in some verification protocols [54, 55], but in our context these settings are further associated with a fidelity estimator defined by the specific measurement basis and outcome.
@@ V.3 Fidelity with W state
The off-diagonal estimator $\hat{o}$ can be found in Appendix A. We have $0\leq\hat{d}\leq 1/n$ and $(n-1)/2\leq\hat{o}\leq(n-1)/2$.
@@
    $\texttt{sum}\,\texttt{+=}\,(n^{2}-n+1)\left(2\delta_{\hat{b}_{i},\hat{b}_{j}}-1\right)\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}/2n$
@@ V.4 Fidelity with Dicke state
We have $0\leq\hat{d}\leq\binom{n}{k}^{-1}$ and $\binom{n}{k}^{-1}\cdot c_{l}\leq\hat{o}_{l}\leq\binom{n}{k}^{-1}\cdot c_{l}$, where the definition of $c_{l}$ and the off-diagonal estimator $\hat{o}_{l}$ can be found in Appendix B. Moreover, let
@@
    $\texttt{sum}\,\texttt{+=}\,2S\binom{n}{k}^{-1}\left(\sum_{|\mathbf{i}|=k}\delta_{\hat{b},\mathbf{i}}-\frac{1}{2}\right)$
@@
        + $|\{m\in\text{s}(\mathbf{j})-\text{s}(\mathbf{i})\mid\hat{b}_{m}=1\}|$
    $\texttt{sum}\,\texttt{+=}\,(-1)^{x}S\cdot\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}/\binom{n}{k}$
@@ VI A generic optimization rule
After restricting the sampling domain to (16) we reweight and renormalize the estimators and the sampling probabilities (which is itself a nontrivial process that depends on both the set (16) and the coefficients $c_{b_{1},b_{2}}$), replacing the original shadow estimation using all $3^{n}$ Pauli operators. An immediate corollary is that Pauli strings containing an odd number of $Y$ are never required when $O\in\mathbb{R}^{2^{n}\times 2^{n}}$.
Suppose the number of nonzero off-diagonal, right triangular elements of $O$ is $r$. Then, according to the measurement rule described above, we obtain a very rough upper bound on the number of measurement settings required to estimate $\text{tr}(\rho O)$, which is $1+rd/2$. If $r=\text{poly}(n)$, this already yields an exponential reduction compared to the total number of local Pauli measurement settings, which is $3^{n}=\left(\frac{3}{2}\right)^{n}d$.
@@
In quantum machine learning, one must estimate the quantity
$\text{tr}(|\psi(\mathbf{x};\bm{\theta})\rangle\langle\psi(\mathbf{x};\bm{\theta})|O),$
where
$|\psi(\mathbf{x};\bm{\theta})\rangle=\prod_{l=1}^{L}V^{(l)}(\bm{\theta})U^{(l)}(\mathbf{x})|0\rangle^{\otimes n}$
for parameterized unitaries $V^{(l)}(\bm{\theta})$ and data-encoding unitaries $U^{(l)}(\mathbf{x})$. There is growing interest in applying shadow tomography to observable estimation in variational quantum neural networks [59, 60]. As our approach can target a wide range of observables, it naturally extends to such applications.
@@ Appendix A Off-diagonal terms for W state
where $n]\equiv\{1,2,\cdots,n\}$. The third equality follows from previous analyses to reduce the sample space.
From Table 1, we see that the term inside the square bracket in (A) is zero if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=0$, $\pm 1/4$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}=U_{j}$, and $\pm i/4$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}\neq U_{j}$. Meanwhile, $\text{tr}(\rho|e_{j}\rangle\langle e_{i}|)$ is the complex conjugate of (A), and the sample
@@
evaluates to $\pm 1/2$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}=U_{j}$. Otherwise, it is always zero. This symmetry allows us to further reduce the sample space of $(U_{i})$:
@@
where $\hat{o}\equiv(n-1)V_{\text{W}}$. Therefore, we have $(n-1)/2\leq\hat{o}\leq(n-1)/2$.
@@ Appendix B Off-diagonal terms for Dicke state
[…]
@@
From Table 1, we see that the term inside the square bracket in (B) is zero if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}=0$ or $\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=0$, $\pm 1/2^{2k-2l}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and $|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 0\pmod{2}$, and $\pm i/2^{2k-2l}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and $|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 1\pmod{2}$. Meanwhile, $\text{tr}(\rho|\mathbf{j}\rangle\langle\mathbf{i}|)$ is the complex conjugate of (B), and the sample
@@
evaluates to $\pm 1/2^{2k-2l-1}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and $|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 0\pmod{2}$. Otherwise, it is always zero. This symmetry allows us to further reduce the sample space of $(U_{m})$:
@@
[…]
where $\hat{o}_{l}\equiv V_{\text{D}}\cdot\binom{n}{k}^{-1}\cdot c_{l}\cdot 2^{2k-2l-1}$ and $\mathcal{F}(\{U_{m}^{\prime}\})\in T\setminus\{\mathbb{I}\}$ such that $|\{U\in\{U_{m}^{\prime}\}\cup\mathcal{F}(\{U_{m}^{\prime}\})|U=HS^{\dagger}\}|\equiv 0\pmod{2}$. Therefore, we have $\binom{n}{k}^{-1}\cdot c_{l}\leq\hat{o}_{l}\leq\binom{n}{k}^{-1}\cdot c_{l}$.